HUBBLE SPACE TELESCOPE DAILY REPORT #841 PERIOD COVERED: 0800 EST Daily Status Report as of 070/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 3538 ( UV spectra of Extended Emission Regions Around QSOs) The Faint Object Spectrograph (Blue Side) was used to make a set of four observations of quasi-stellar object B2-1512+37 to obtain spectra of the extended emission line region around the quasar. All four observations were received and displayed. The first was the binary target acquisition (eight groups), the second was the confirmation image, and the last two were the scientific spectral observations. The target was well centered in the 4.3 arcsecond aperture. The spectral observations were reported to be weak and noisy, but showed some signs of emission lines at the expected wave lengths. The third observations was filed as being bad because of high jitter during a night/day transition. 1.2 Completed Two Sets of S/C/FOS/BL 4099 (FOS focus, X-Pitch and Y-Pitch) The Faint Object Spectrograph was used to make two sets (of seven observations each) as part of the focus and x/y pitch test. All fourteen observations were received and displayed. They were reported to be okay. 1.3 Completed FOC/96 4800 (Stellar content of Wolf-Rayet Galaxies) The Faint Object Camera (f/96) was used to make one image of the Wolf-Rayet Galaxy NGC-1741. The star burst galaxy is know to exhibit broad He II emission lines from stellar sources. The observations was executed without problems. The data was received and displayed. There was roll-over in the central pixel for some of the brightest objects, and the data looked good. 1.4 Completed WFC 4658 (Determination of Extragalactic Distance Scale) An attempt was made to use the Wide Field Camera to make observations of NGC-5457 (aka, M-101, a large spiral galaxy in Ursa Major). The baseline guide star acquisition failed , because the FHST update preceding it corrected only for roll, and not for v2-v3 error. The subsequent reacquisition was executed with out problems and it was received and was displayed. The first image was empty, and the second was filled with stars and cosmic ray hits. 1.5 Completed GHRS/2 3626 (Ephemeral Determination fo wind Velocities and Density Laws) The High Resolution Spectrograph was used to make a series of eight observations of the star Zeta Aurigae (aka, HD32068). The onboard spiral search, and the peakup were successful. The wave length calibrations, and the Echelle-B spectral observations were executed without problems. 1.6 Completed PC 4140 (WF/PC Calibration: Rapid Internal Monitor Part 2) The Planetary Camera was used to make internal flat observations to check for new forms of contamination, and the reappearance of quantum efficiency hysteresis (QEH). The observations were executed as scheduled, and no problems were reported. 1.7 Completed Uplink Updated Ephemeris Tables, Per MOSES 2871 1.8 Completed GHRS/2 3496 (Ultraviolet Atlas of Sirius) The High Resolution spectrograph was used to obtain ultraviolet spectra of the star Sirius (aka, HD48915) in the WSCAN Mode to get an accurate measurements of its flux distribution and line profiles. The observations were executed as scheduled. There was an acquisition failure at 070/0807Z and a reacquisition failure at 070/0943Z. The observations were filed as bad. 1.9 Initiated FOC/96 3591 (Massive Stars in Starburst galaxies - FOC Images) The Faint Object Camera was used to initiate ultraviolet images of the star bursting galaxy IZW18. The observations are currently under way, and will be analyzed later. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 3 (Failure at 069/2104Z, Attitude error, reason is two V1-Axis updates after a major slew.) (Failure 070/0807Z, the attitude looked good, but no primary star found which exceeded the threshold.) Scheduled Reacquisitions : 7 Successful : 6 (Failure at 070/0943) Losses of Lock : 0 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Significant HSTARs: None 2.4 Operations Note: The High Speed Photometer was successfully recovered from Safe Mode yesterday (Wednesday 069/1836Z) using real-time commands. No problems were reported. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 3591 (Massive Stars in Starburst Galaxies - FOC Images) 3.2 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.3 Complete FOC/96 4800 (Stellar Content of Wolf-Rayet Galaxies) 3.4 Complete HSP/POL Polorimetric Observations of the Crab Nebula Pulsar in the UV) 3.5 Complete WFC 1306 (GTO Low Latitude Parallel Program) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-